User:Jason O Archer/Notebook/PHYC 307L Junior Lab/2008/09/15

From OpenWetWare

Jump to: navigation, search
PHYC 307L Junior Lab Main project page
Previous entry      Next entry

Planck's constant lab 9/15/08

After obtaining the mercury lamp, voltmeter, and specialized h/e apparatus, I had some confusion regarding a conflict in the lab manual and the apparatus; the manual indicated warming up the lamp for five minutes, but the mercury lamp itself indicated warming up for twenty minutes. On Dr. Koch's suggestion, I waited for twenty minutes, in order to allow the mercury lamp to stabilize.

Upon checking the battery voltages on the two blue battery terminals, I discovered them to be between 1 and 2 volts, far from the expected 6 V. I tried getting new cables for the voltmeter, and the low voltage persisted. I tried connecting to different terminals, and the low voltages persisted. I noticed that the magnetized diffraction gratings were attached to the apparatus; I removed them, and the low voltages persisted. I tried pressing the "PUSH TO ZERO" button a few times, in case there was any residual charge affecting the readings; the low voltages persisted. At this point I tried checking the batteries--they were dead. With Aram's help, I replaced the batteries; he tested the new batteries to make sure they were functional (they were).

After replacing the batteries, I measured the voltages again; the negative voltage was around -9.4 V, and the positive voltage was around +9.4 V; these figures were well within acceptable minima of -6 V and +6 V, respectively. SJK 12:32, 20 October 2008 (EDT)
12:32, 20 October 2008 (EDT)good thing you checked this, and persisted...thank you for replacing the batteries!
12:32, 20 October 2008 (EDT)
good thing you checked this, and persisted...thank you for replacing the batteries!

I tried to focus the yellow/green wavelengths of light from the lamp on the aperture window of the h/e apparatus. While doing this, I noticed a possible source of error; the screw to hold the aperture in place is missing. Aram advised me that this should not be an issue in determining data, as long as I do not move the aparratus.

I turned on the apparatus and focused on the green maximum in the first order on the right; I pushed the "PUSH TO ZERO" button until the voltmeter read no voltage, then released it. After several seconds, the initial reading fluctuated between 1.237 V and 1.239 V.

Photon Theory of Light, Part A

From the same green line shown above, with the green filter and the 100% transmission filter added, the initial reading of stopping potential fluctuated between 0.886 V and 0.887 V. on the first reset, it returned to the previously fluctuating voltage. On the second reset, it took 54.38 seconds to return to that voltage. On the third reset, it took 52.68 seconds. On the 4th reset, it took 44.62 s. On the 5th reset, it took 49.41 s. On the 6th, it took 39.91 s.

From the green line, with the green filter and the 80% transmission filter, the initial reading of stopping potential fluctuated between 0.886 V and 0.887 V. On the first reset, it returned to this voltage in 53.19 s. On the second reset, it took 61.63 s. On the 3rd reset, it took 59.28 s. On the 4th reset, I forgot to reset my stopwatch. On the 5th reset, it took 56.56 s. On the 6th reset, it took 62.53 s.

From the green line, with the green filter and the 60% transmission filter, the initial reading of stopping potential fluctated between 0.885 V and 0.887 V.


Personal tools